# Stochastic Particle Acceleration during Pressure-Anisotropy-Driven Magnetogenesis in the Pre-Structure Universe

### 05 Mar 2026 v1

@@ III.3 Sensitivity to a pre-existing suprathermal population
$f(p,z=10)=f_{\rm MW}(p,z=10)+A\,p^{-s}\exp(-p/p_{\rm cut}).$ (38)
We then evolve this distribution using Eq. (35). The slope of the pre-existing tail is taken to be $s=4.2$, motivated by typical spectra expected from DSA. The normalization of the tail $A$ is chosen such that the nonthermal component matches a small fraction of the thermal Maxwellian at a reference momentum $p_{0}$, while an exponential cutoff $p_{\rm cut}$ limits the high-momentum extent of the seed population.
The resulting spectra are shown in Fig. 5. We find that the overall evolution from $z=10$ to $z=2$ is largely insensitive to the presence of this weak pre-existing suprathermal population. In both turbulence cases ($\epsilon=0.05$ and $\epsilon=0.2$), the proton distribution shifts primarily toward lower momenta as the universe expands. This behavior arises from the adiabatic momentum loss term $\dot{p}_{\rm ad}=-H(z)p$, which implies $p\propto a^{-1}\propto(1+z)$. Although stochastic momentum diffusion slightly broadens the high-momentum tail, particularly for larger turbulence amplitudes, adiabatic cooling remains the dominant influence on the spectral evolution over the interval $z=10\rightarrow 2$.
As the system approaches the magnetic-field turn-on epoch ($z\sim z_{\rm on}\sim 1.7$), the reduction of the stochastic acceleration timescale allows the diffusion term to partially compete with adiabatic losses. Consequently, a mild suprathermal tail begins to emerge in the particle spectrum. The presence of a pre-existing power-law tail does not significantly alter this behavior, indicating that the early evolution is largely controlled by cosmological cooling rather than by stochastic re-acceleration of the seed population.
IV Comparison with shock acceleration in the Bohm-diffusion limit
@@
$t_{\rm acc,sh}(E)\approx\frac{3}{u_{1}-u_{2}}\left(\frac{D_{1}(E)}{u_{1}}+\frac{D_{2}(E)}{u_{2}}\right),$ (39)
where $u_{1,2}$ and $D_{1,2}$ are the upstream/downstream flow speeds and spatial diffusion coefficients, respectively. For a strong shock with compression ratio $r=u_{1}/u_{2}\simeq 4$ and comparable diffusion coefficients on both sides ($D_{1}\sim D_{2}\sim D$), Eq. (39) reduces to the commonly used estimate
$t_{\rm acc,sh}(E)\sim\xi\,\frac{D(E)}{u_{s}^{2}},\qquad\xi\sim\mathcal{O}(10),$ (40)
@@
$D_{\rm B}(E)\equiv\frac{1}{3}r_{L}(E)c,\qquad r_{L}(E)=\frac{pc}{eB}\approx\frac{E}{eB},$ (41)
for relativistic particles with $E\approx pc$. Combining Eqs. (40)–(41) yields
$t_{\rm acc,sh}(E)\sim\frac{\xi}{3}\,\frac{r_{L}(E)c}{u_{s}^{2}}\sim\frac{\xi}{3}\,\frac{E}{eB}\,\frac{c}{u_{s}^{2}}.$ (42)
Timescale ratio and dominance of shocks.
@@
$\frac{t_{\rm acc,2}}{t_{\rm acc,sh}}\sim\frac{3}{\xi}\,\left(\frac{c}{v_{\rm tur}}\right)^{2}\left(\frac{u_{s}}{c}\right)^{2}\left(\frac{c}{r_{L}\,\nu_{\rm eff}}\right),$ (43)
where $\xi\sim\mathcal{O}(10)$ parameterizes order-unity factors in the DSA timescale. The final factor, $c/(r_{L}\nu_{\rm eff})$, measures the efficiency of pitch-angle scattering over a gyro-orbit; in the Bohm-like limit, $\nu_{\rm eff}\sim c/r_{L}$, this factor is of order unity.
For structure-formation environments, the relevant characteristic velocities are $v_{\rm tur}\sim\epsilon c_{s}$ with $c_{s}\sim 10^{2}-10^{3}~{\rm km\,s^{-1}}$ and $\epsilon\sim 0.1$, implying $v_{\rm tur}/c\sim 10^{-4}-10^{-3}$. Large-scale structure shocks typically have $u_{s}\sim 10^{3}~{\rm km\,s^{-1}}$, corresponding to $u_{s}/c\sim 3\times 10^{-3}$. Adopting Bohm-like scattering ($c/(r_{L}\nu_{\rm eff})\sim 1$), these scalings imply
$\frac{t_{\rm acc,2}}{t_{\rm acc,sh}}\sim 10-10^{3},$ (44)
@@
V Summary and Discussion
We investigated whether stochastic (second-order Fermi) acceleration during pressure-anisotropy-driven magnetogenesis can generate a dynamically significant CR population prior to nonlinear structure formation. By combining analytic acceleration-time constraints with a Fokker-Planck model, we quantified both the turn-on condition for efficient stochastic acceleration and the resulting proton momentum distribution in the early Universe. The analytic criterion $t_{\rm acc}<t_{H}$ yields a critical magnetic field $B_{\rm crit}(z)$ and defines a CR turn-on redshift $z_{\rm on}\sim 1.7$ for representative parameters. At higher redshift, the magnetic field remains below this threshold, so that stochastic acceleration proceeds too slowly to compete with cosmological expansion. Even near the turn-on epoch, the maximum attainable proton energy, evaluated under optimistic assumptions corresponding to the strong-scattering limit, reaches at most $\mathcal{O}(10^{2})\,\mathrm{GeV}$. The numerical Fokker-Planck calculation shows that, throughout $z=10\rightarrow z_{\rm on}$, the proton distribution remains close to a cooling Maxwellian. Adiabatic expansion dominates the evolution, and the resulting nonthermal energy density remains far below the thermal background energy density.
These results carry several cosmological implications. First, instability-mediated stochastic acceleration is unlikely to modify the thermal history of the pre-structure intergalactic medium. Second, the resulting suprathermal component is too weak to produce observable distortions of the cosmic microwave background or to contribute significantly to diffuse gamma-ray or neutrino backgrounds. Finally, efficient CR production appears to be intrinsically tied to the onset of nonlinear structure formation, when large-scale shocks and amplified magnetic fields operate on substantially shorter acceleration timescales. While the low-level suprathermal population found here may in principle serve as a seed for subsequent DSA, its energy content is modest. A detailed assessment of the injection efficiency and the reprocessing of this seed population by structure-formation shocks will require a dedicated kinetic calculation, which we defer to future work.
In summary, pressure-anisotropy-driven magnetogenesis can modulate particle scattering in the early Universe, but it does not by itself generate a dynamically important pre-shock CR population. The emergence of efficient CR acceleration is therefore naturally associated with the epoch of nonlinear structure formation.

### 01 Apr 2026 v2

@@ 4.3 Limited impact of a pre-existing suprathermal population
$f(p,z=10)=f_{\rm MW}(p,z=10)+Ap^{-s}\exp(-p/p_{\rm cut}).$ (4.11)
The normalization of the tail is fixed by matching it to the Maxwellian distribution at a reference momentum $p_{0}$, such that
$A,p_{0}^{-s}=\eta_{\rm tail}f_{\rm MW}(p_{0},z=10),$ (4.12)
where $\eta_{\rm tail}\ll 1$ controls the relative amplitude of the suprathermal component. This condition ensures a smooth connection between the thermal and nonthermal populations at $p=p_{0}$. We then evolve this distribution using Eq. (4.8). The slope of the pre-existing tail is taken to be $s=4.2$, representative of typical nonthermal spectra in collisionless plasmas, and consistent with expectations from both stochastic acceleration and shock-related processes. An exponential cutoff $p_{\rm cut}$ limits the high-momentum extent of the seed population.
Figure 6: Same as Figure 5, but including a weak pre-existing suprathermal power-law tail ($s=4.2$) in the initial distribution.
The resulting spectra are shown in Figure 6. We find that the overall evolution from $z=10$ to $z=2$ is largely insensitive to the presence of a weak pre-existing suprathermal population. In both panels, the ion distribution shifts toward lower momenta as the universe expands, reflecting adiabatic momentum losses $\dot{p}_{\rm ad}=-H(z)p$, for which $p\propto a^{-1}\propto(1+z)$. However, when Coulomb losses are included, a clear departure from purely adiabatic evolution appears at low momenta. In particular, particles with $p/(m_{i}c)\lesssim 5\times 10^{-3}$, corresponding to kinetic energies of order $\sim 10~{\rm keV}$, are efficiently thermalized by Coulomb interactions. In this regime, the cooling timescale is shorter than the acceleration timescale, and stochastic diffusion is effectively suppressed. As a result, the pre-existing power-law tail does not survive at low energies, and the distribution remains close to a Maxwellian core throughout the evolution. At higher momenta ($p/(m_{i}c)\gtrsim 5\times 10^{-3}$), the Coulomb loss timescale increases rapidly, allowing stochastic acceleration to operate. Nevertheless, even in this regime, the resulting modification of the spectrum remains modest over the interval $z=10\rightarrow 2$, indicating that cosmological cooling continues to dominate the overall evolution.
As the system approaches the magnetic-field turn-on epoch ($z\sim z_{\rm on}\sim 1.7$), the reduction of the stochastic acceleration timescale allows the diffusion term to partially compete with losses. However, the results indicate that Coulomb cooling still limits the efficient re-acceleration of the seed population. The presence of a pre-existing tail therefore does not significantly enhance the formation of a nonthermal component in the pre-shock phase. Instead, the early evolution is controlled primarily by the competition between Coulomb thermalization and cosmological expansion, with stochastic acceleration becoming effective only above the effective threshold set by Coulomb losses.
5 Comparison with shock acceleration in the Bohm-diffusion limit
@@
$t_{\rm acc,sh}(E)\approx\frac{3}{u_{1}-u_{2}}\left(\frac{D_{1}(E)}{u_{1}}+\frac{D_{2}(E)}{u_{2}}\right),$ (5.1)
where $u_{1,2}$ and $D_{1,2}$ are the upstream/downstream flow speeds and spatial diffusion coefficients, respectively. For a strong shock with compression ratio $r=u_{1}/u_{2}\simeq 4$ and comparable diffusion coefficients on both sides ($D_{1}\sim D_{2}\sim D$), Eq. (5.1) reduces to the commonly used estimate
$t_{\rm acc,sh}(E)\sim\xi\,\frac{D(E)}{u_{s}^{2}},\qquad\xi\sim\mathcal{O}(10),$ (5.2)
@@
$D_{\rm B}(E)\equiv\frac{1}{3}r_{L}(E)c,\qquad r_{L}(E)=\frac{pc}{eB}\approx\frac{E}{eB},$ (5.3)
for relativistic particles with $E\approx pc$. Combining Eqs. (5.2)–(5.3) yields
$t_{\rm acc,sh}(E)\sim\frac{\xi}{3}\,\frac{r_{L}(E)c}{u_{s}^{2}}\sim\frac{\xi}{3}\,\frac{E}{eB}\,\frac{c}{u_{s}^{2}}.$ (5.4)
@@
$\frac{t_{\rm acc,2}}{t_{\rm acc,sh}}\sim\frac{3}{\xi}\,\left(\frac{c}{v_{\rm tur}}\right)^{2}\left(\frac{u_{s}}{c}\right)^{2}\left(\frac{c}{r_{L}\,\nu_{\rm eff,B}}\right),$ (5.5)
where $\xi\sim\mathcal{O}(10)$ parameterizes order-unity factors in the DSA timescale. The final factor, $c/(r_{L}\nu_{\rm eff,B})$, measures the efficiency of pitch-angle scattering over a gyro-orbit; in the Bohm-like limit, $\nu_{\rm eff,B}\sim c/r_{L}$, this factor is of order unity.
For structure-formation environments, the relevant characteristic velocities are $v_{\rm tur}\sim\epsilon c_{s}$ with $c_{s}\sim 10^{2}-10^{3}~{\rm km\,s^{-1}}$ and $\epsilon\sim 0.1$, implying $v_{\rm tur}/c\sim 10^{-4}-10^{-3}$. Large-scale structure shocks typically have $u_{s}\sim 10^{3}~{\rm km\,s^{-1}}$, corresponding to $u_{s}/c\sim 3\times 10^{-3}$. Adopting Bohm-like scattering ($c/(r_{L}\nu_{\rm eff,B})\sim 1$), these scalings imply
$\frac{t_{\rm acc,2}}{t_{\rm acc,sh}}\sim 10-10^{3},$ (5.6)
@@
6 Summary and Discussion
We have investigated whether stochastic (second-order Fermi) acceleration associated with pressure-anisotropy-driven magnetogenesis can generate a dynamically significant population of cosmic rays (CRs) prior to the onset of nonlinear structure formation. By combining analytic acceleration-time constraints with a Fokker–Planck description of particle transport, we quantified both the conditions under which stochastic acceleration becomes cosmologically viable and the resulting ion momentum distribution in the pre-structure intergalactic medium.
An analytic comparison between the acceleration timescale and the Hubble time yields a critical magnetic field $B_{\rm crit}(z)$ and defines a characteristic turn-on redshift $z_{\rm on}\sim 1.7$, at which stochastic acceleration can in principle begin to compete with cosmological expansion. We find that this turn-on epoch is largely insensitive to uncertainties in the thermal history and microphysical parameters, reflecting the rapid nonlinear amplification of the magnetic field. However, this condition provides only an optimistic upper bound on the efficiency of particle acceleration.
When additional energy loss processes are included, the situation changes qualitatively. In particular, Coulomb interactions in a fully ionized intergalactic medium introduce a strong energy dependence in the effective loss timescale, leading to a refined critical magnetic field $B_{\rm crit}^{\rm eff}(E,z)$. This analysis reveals the existence of an effective injection threshold for stochastic acceleration at energies of order ${\cal O}(10)$ keV, corresponding to $p/(m_{i}c)\sim 5\times 10^{-3}$. Below this scale, Coulomb losses dominate over acceleration ($t_{\rm C}\ll t_{\rm acc}$), efficiently thermalizing particles and suppressing their injection into the acceleration process. At higher energies, where $t_{\rm C}\gg t_{H}$, the acceleration condition approaches the Hubble-limited regime, and the maximum attainable energy remains essentially unchanged from the adiabatic-only estimate.
To quantify the resulting particle population, we solved a Fokker–Planck equation for the isotropic ion distribution over the redshift interval $z=10\rightarrow z_{\rm on}$. The numerical results show that, even under optimistic assumptions corresponding to the strong-scattering limit, the ion distribution remains close to a cooling Maxwellian throughout this epoch. Adiabatic expansion dominates the evolution, and Coulomb losses further suppress the formation of a suprathermal tail by efficiently removing particles near the injection scale. Even when a pre-existing suprathermal component is included in the initial condition, it is rapidly thermalized at low energies and does not lead to a significant enhancement of the nonthermal population.
These results imply that stochastic acceleration during pressure-anisotropy-driven magnetogenesis is unable to generate a dynamically important CR population in the pre-structure universe. In particular, Coulomb losses transform the acceleration problem from a purely cosmological constraint into an energy-dependent injection problem, strongly limiting the fraction of particles that can participate in stochastic energization. As a consequence, the resulting suprathermal population remains energetically subdominant and is unlikely to have a measurable impact on the thermal history of the intergalactic medium or on observable high-energy backgrounds.
Our findings therefore support a picture in which efficient CR production is intrinsically linked to the onset of nonlinear structure formation. While instability-driven scattering can enhance particle transport and may provide a modest level of pre-acceleration, the emergence of a substantial nonthermal population requires the presence of large-scale shocks and sufficiently amplified magnetic fields. The low-level suprathermal population obtained here may serve as a seed for subsequent diffusive shock acceleration (DSA), but its contribution to the overall CR energy budget is expected to be limited. A detailed kinetic study of particle injection and re-acceleration at structure-formation shocks, incorporating the effects of pre-existing turbulence and seed populations, is left for future work.
Acknowledgments
This work was carried out with the support of the Korea Astronomy and Space Science Institute (KASI). The author thanks the referees for their careful reading of the manuscript and for their constructive comments and suggestions, which have significantly improved the quality and clarity of this work.